# V2P: From Background Suppression to Center Peaking for Robust GUI Grounding Task

### v1

@@ 4.2 Main Result
Table 1: Comparison of Model Performance Across Task Categories in ScreenSpot-Pro. Bold text highlights the best results, while “–” represents missing values not reported in the original papers. Details of selected baselines are in Appendix C
Model ScreenSpot-v2 Accuracy (%)
Proprietary Models
Operator 47.3 41.5 90.2 80.3 92.8 84.3 70.5
GPT-4o + OmniParser-v2 95.5 74.6 92.3 60.9 88.0 59.6 80.7
General Open-source Models
Qwen2.5-VL-3B 93.4 73.5 88.1 58.6 88.0 71.4 80.9
Qwen2.5-VL-7B 97.6 87.2 90.2 74.2 93.2 81.3 88.8
GUI-specific Models (SFT)
SeeClick-9.6B 78.4 50.7 70.1 29.3 55.2 32.5 55.1
Magma-8B 62.8 53.4 80.0 57.9 67.5 47.3 61.5
OS-Atlas-4B 87.2 59.7 72.7 46.4 85.9 63.1 71.9
UI-TARS-2B 95.2 79.1 90.7 68.6 87.2 78.3 84.7
OS-Atlas-7B 95.2 75.8 90.7 63.6 90.6 77.3 84.1
Aguvis-7B 95.5 77.3 95.4 77.9 91.0 72.4 86.0
UGround-V1-7B 95.0 83.3 95.0 77.8 92.1 77.2 87.6
UI-TARS-72B 94.8 86.3 91.2 87.9 91.5 87.7 90.3
GUI-Actor-3B 97.6 83.4 96.9 83.6 94.0 85.7 91.0
UI-TARS-7B 96.9 89.1 95.4 85.0 93.6 85.2 91.6
GUI-Actor-7B 97.6 88.2 96.9 85.7 93.2 86.7 92.1
GUI-specific Models (RL)
SE-GUI-7B - - - - - - 90.3
LPO-8B - - - - - - 90.5
Ours
V2P-7B 98.1 88.0 96.1 89.7 95.4 84.4 92.3
4.3 Ablation Study
4.3.1 Ablation Study for V2P
The ablation study in Tab. 3 highlights the effectiveness of our V2P method. On the challenging ScreenSpot-Pro, removing Fitts-Gaussian Labeling results in a drop of 3.0% (50.5% to 47.5%), showing its importance in resolving center-edge confusion. Further removing Suppression Attention leads to an additional decline of 3.2%, underscoring its role in reducing background distractions. These results illustrate that first filtering out irrelevant regions to isolate target areas, followed by refining localization to pinpoint the center of UI elements with precision offering clear advantages in addressing complex scenarios.
On the simpler dataset ScreenSpot-v2, performance remains nearly unchanged after removing Fitts-Gaussian Labeling (92.3%), as the dataset’s straightforward tasks already allow high accuracy. And we suspect this is also because ScreenSpot-v2 primarily features simple UI layouts with minimal overlap, as Fitts-Gaussian Labeling is especially effective on more complex interfaces with overlapping UI elements, which guides the model to click precisely at the icon center, reducing misclicks on adjacent elements (we demonstrate it in Sec. 4.3.3). This indicates that while V2P provides limited gains on easier datasets like ScreenSpot-v2, its full potential shines in challenging real-world settings like ScreenSpot-Pro. However, removing both components results in a slight drop (91.9%), particularly in the Mobile-Icon scenario.
ScreenSpot-Pro
Model Dev Creative CAD Scientific Office OS Avg.
V2P-7B (Full) 46.8 43.1 47.1 56.3 68.3 45.4 50.5
   w/o Fitts-Gaussian Labeling 42.8 39.9 44.0 49.2 67.8 46.4 47.5 (-3.0 $\downarrow$)
   w/o Fitts-Gaussian & Suppression Attention 43.5 41.1 38.3 49.6 63.5 43.4 44.3 (-6.2 $\downarrow$)
ScreenSpot-v2
Model Mobile-Text Mobile-Icon Desktop-Text Desktop-Icon Web-Text Web-Icon Avg.
V2P-7B (Full) 98.1 88.0 96.11 89.7 95.4 84.4 92.3
   w/o Fitts-Gaussian Labeling 98.8 86.4 97.2 87.3 95.4 85.7 92.3 (-0.0)
   w/o Fitts-Gaussian & Suppression Attention 98.1 83.3 95.6 88.1 95.7 86.9 91.9 (-0.4 $\downarrow$)
Table 3: Ablation study on ScreenSpot-Pro and ScreenSpot-v2.
Figure 4: Ablation studies of Fitts-Gaussian Modeling (FGPM). (a) Effect of the Gaussian factor $\sigma$ on overall accuracy for ScreenSpot-v2 and ScreenSpot-Pro. (b) Size-wise performance on ScreenSpot-Pro.
4.3.2 Ablation Study for Gaussian Factor $\sigma$
We conducted ablation experiments to analyze the effect of different Gaussian factors $\sigma$ on model performance. As shown in Fig. 4(a), the model’s performance is strongly influenced by the choice of Gaussian factor $\sigma$. For both ScreenSpot-v2 and ScreenSpot-Pro, accuracy improves as $\sigma$ decreases. For example, on ScreenSpot-v2, the accuracy rises from 91.3% at $\sigma=6.0$ to 92.4% at $\sigma=0.5$, while ScreenSpot-Pro achieves its best result of 50.5% accuracy at $\sigma=1.0$.
We suspect that this is because that larger $\sigma$ values correspond to a broader Gaussian distribution, which tends to dilute the spatial focus and introduce noise into the attention maps. In comparison, smaller $\sigma$ produces sharper Gaussian peaks, allowing the model to localize UI elements with higher precision and resulting in more accurate click predictions. These results underscore that carefully balancing the Gaussian factor is crucial: excessively large values hinder localization, while moderate to smaller values significantly enhance spatial accuracy and overall model performance.
4.3.3 Ablation Study for Effectiveness of Fitts-Gaussian Peak Modeling
Traditional attention methods often yield overly broad regions, misaligning with small UI elements and producing points outside their boxes (Fig. 2(b)). Fitts-Gaussian Peak Modeling counters this by centering the attention, boosting accuracy on tiny elements. We ablate the design on the challenging ScreenSpot-Pro dataset Li et al. (2025), splitting its boxes into small, medium, and large and comparing results with/without our Fitts-Gaussian Peak Modeling (FGPM).
The results in Fig. 4(b) show that the proposed modeling yields substantial performance improvements for small UI elements by 7.4%, and also provides noticeable gains for medium elements (3.0%). However, for large UI elements, we observe a slight decrease in performance. We attribute this to the fact that our method’s centralized attention may sometimes neglect the full extent of the target, whereas the broader attention maps in prior methods might coincidentally localize such large elements better. Nevertheless, our approach still brings substantial overall improvement, especially in more challenging cases.
5 Conclusion
We propose Valley-to-Peak (V2P), a framework that addresses key challenges in leveraging vision-language model attentions for GUI grounding tasks (i.e., background distraction and center-edge confusion). By introducing Suppression Attention to reduce distractions and Fitts-Gaussian Labeling to guide the model toward actionable centers, V2P enhances the model’s ability to focus precisely on target elements, better simulating human interaction behavior. Experiments show V2P achieves modest gains on simpler datasets like ScreenSpot-v2 (92.3%, +3.5%) but excels in complex settings like ScreenSpot-Pro (50.5%, +23.7%). Ablation studies confirm the effectiveness of both components, highlighting V2P’s robustness and applicability to high-precision GUI grounding tasks.
Appendix A Training Details
A.1 Source Training Data
Following GUI-Actor Wu et al. (2025), we compile our training dataset from several publicly available, high-quality GUI datasets, with summary statistics provided in Table 4. To ensure fair evaluation, we also exclude any samples from Wave-UI that overlap with the test sets of downstream tasks.
Dataset # of Elements # of Screenshots Platform
Uground Web–Hybrid Gou et al. (2025a) 8M 775K Web
GUI-Env Chen et al. (2025) 262K 70K Web
GUI-Act Chen et al. (2025) 42K 13K Web
AndroidControl Li et al. (2024) 47K 47K Android
AMEX Chai et al. (2025) 1.2M 100K Android
Wave-UI 50K 7K Hybrid
Total 9.6M 1M
Table 4: Overview of training datasets used for GUI-Actor.
A.2 Training and Inference Details
During the training phase, we first freeze the backbone VLM parameters and train only the action head ($\sim 20$M parameters). In the second phase, we fine-tune the entire model using the filtered dataset with standard supervised learning. At inference, we follow deterministic generation with a temperature of 0 and adopt a confidence threshold of $\gamma=0.95$ for the ScreenSpot-Pro benchmark and $\gamma=0.8$ for ScreenSpot-v2 tasks.
Appendix B Benchmarks
Our evaluation centers on two sophisticated benchmarks for GUI visual grounding: ScreenSpot-v2 Wu et al. (2024b) and ScreenSpot-Pro Li et al. (2025).
ScreenSpot-v2 encompasses 1,272 carefully annotated instructions, each paired with corresponding target elements across diverse GUI environments, including mobile (Android and iOS), desktop (macOS and Windows), and web platforms. The dataset is designed to improve the quality and reliability of GUI visual grounding tasks, addressing key challenges such as eliminating ambiguities in natural language instructions and resolving annotation errors. By refining the alignment between textual descriptions and interface elements, ScreenSpot-v2 provides a robust and standardized benchmark for evaluating grounding models.
ScreenSpot-Pro, meanwhile, focuses on more demanding scenarios, especially those involving high-resolution professional applications. It contains 1,581 tasks annotated by domain experts across 23 specialized software applications, spanning three operating systems. This benchmark significantly broadens the scope of GUI visual grounding by introducing interfaces with industrial software and multi-window layouts, creating a larger domain gap compared to most pretraining data. With its increased complexity and domain diversity, ScreenSpot-Pro is an invaluable resource for assessing the generalization ability of models in realistic and challenging GUI environments.
Appendix C Baselines
C.1 Baselines for ScreenSpot-Pro
We establish comprehensive benchmarking across four categories of state-of-the-art GUI understanding systems:
  • Proprietary Systems: GPT-4o OpenAI (2024) (vision-language foundation model), Claude Computer Use Google (2024) (specialized GUI agent)
  • General-Purpose Open-Source: Qwen2.5-VL series Bai et al. (2025) (7B/72B parameter variants)
  • GUI-Specialized (SFT):
    • Medium-scale: SeeClick-9.6B Cheng et al. (2024), FOCUS-2B Zhang et al. (2024), OS-Atlas-7B Wu et al. (2024a)
    • Large-scale: CogAgent-18B Hong et al. (2024), Aria-UI Yang et al. (2025b), JEDI series Xie et al. (2025)
    • Domain-specific: ShowUI-2B Lin et al. (2024), Uground series Gou et al. (2025b), UI-TARS series Qin et al. (2025)
  • GUI-Specialized (RL):
    • R1-style: UI-R1 Lu et al. (2025), GUI-R1 Luo et al. (2025), InfiGUI-R1-3B Liu et al. (2025)
    • Gaussian-based: GUI-G1-3B Zhou et al. (2025), SE-GUI Yuan et al. (2025), GUI-$\text{G}^{2}$-7B Tang et al. (2025a)
C.2 Baselines for ScreenSpot-v2
We establish comprehensive benchmarking across four categories of state-of-the-art GUI understanding systems:
  • Proprietary Systems: Operator OpenAI (2023) (proprietary multimodal system), GPT-4o + OmniParser-v2 OpenAI (2024); Wan et al. (2024) (enhanced vision-language integration)
  • General-Purpose Open-Source: Qwen2.5-VL series Bai et al. (2025) (7B/72B parameter variants)
  • GUI-Specialized (SFT):
    • Medium-scale: SeeClick-9.6B Cheng et al. (2024), Magma-8B Yang et al. (2025a), OS-Atlas series Wu et al. (2024a)
    • Large-scale: UI-TARS series Qin et al. (2025), Uground series Gou et al. (2025b), GUI-Actor series Wu et al. (2025)
    • Domain-specific: Aguvis-7B Xu et al. (2025)
  • GUI-Specialized (RL):
    • SE-GUI-based: SE-GUI-7B Yuan et al. (2025), LPO-8B Tang et al. (2025b)

### v3

@@ Appendix C Detailed Experimental Results on ScreenSpot-v2 and ScreenSpot-Pro
Table 7: Comparison of Model Performance Across Task Categories in ScreenSpot-Pro. Bold text highlights the best results, while “–” represents missing values not reported in the original papers. The baseline models utilize various backbones and parameter sizes, as indicated by their names (e.g., -7B, -18B).
Appendix D Qualitative Analysis and Case Studies
D.1 Success Cases
Fig. 4 demonstrate several representative success cases where our V2P-7B model achieves accurate GUI element localization. Through these successful examples, we observe that the model exhibits high confidence in precisely highlighting target regions, with attention distributions that closely align with the actual shapes of UI elements. The attention maps show sharp, well-defined boundaries that accurately correspond to button edges, text field borders, and icon contours. This demonstrates the model’s robust understanding of visual-semantic correspondence between natural language instructions and GUI components, effectively bridging the gap between textual descriptions and visual interface elements.
D.2 Failure Cases and Error Analysis
Our analysis of failure cases reveals several interesting patterns and limitations, as illustrated in Fig. 5. In some instances, we observe that the model encounters difficulties when multiple UI elements share semantic similarities. The model often exhibits high confidence while incorrectly selecting semantically related but functionally different elements or misidentifying similar icons with different purposes (Fig. 5(a)).
Additionally, we identify cases where the model’s attention distribution becomes highly dispersed across the interface, which we interpret as an indicator of low confidence (Fig. 5(b)). This scattered attention pattern typically occurs in scenarios with numerous distracting elements or cluttered interfaces, suggesting that the model’s decision-making process becomes uncertain when faced with complex visual layouts.
Furthermore, we observe failure modes where the model’s attention concentrates entirely on regions completely unrelated to the target element (Fig. 5(c)). These cases often involve ambiguous natural language descriptions or interfaces with unconventional design patterns that deviate from the model’s training distribution. Such failures highlight the need for enhanced user intent understanding and more comprehensive UI context comprehension capabilities.
D.3 Multi-step Interaction Scenarios
To visualize the model’s capability in maintaining context across sequential operations, we present case studies of multi-step workflows from the AndroidControl (Li et al., 2024) dataset. Fig. 6(a) and Fig. 7 showcases the model’s performance across sequential GUI operations.
The results demonstrate that our model maintains consistent accuracy throughout extended interaction sequences, successfully completing multi-step tasks that require contextual understanding and state awareness.
D.4 Multi-target Localization Capabilities
We investigated the model’s ability to simultaneously localize multiple targets within a single interface, which holds significant value for batch operations and improving inference efficiency. Fig. 6(b) presents our experimental setup using a calculator interface, where we tasked the model with simultaneously localizing the elements "1", "0", and "00".
The results reveal that the model successfully generates attention distributions for all three target elements simultaneously, with appropriately differentiated confidence levels. Notably, the element "1" receives the highest attention intensity, followed by "0" and "00" respectively, which aligns with the natural priority of these elements. This multi-target capability demonstrates the model’s sophisticated attention mechanism and its potential for complex GUI analysis tasks requiring simultaneous element identification, as well as its genuine understanding capability of user queries.
(a) Success Case 1
(b) Success Case 2
(c) Success Case 3
Figure 4: Representative success cases of GUI element localization.
(a) Failure Case 1
(b) Failure Case 2
(c) Failure Case 3
Figure 5: Representative failure cases of GUI element localization.
(a) Multi step grounding case 1: "Open Phase of the moon App, select the date 25 July on the calendar and view the moon phase for that date." Step 1 (left) and Step 2 (right).
(b) Multi-target grounding case.
Figure 6: Multi-step grounding case and multi-target grounding case.
(a) Step 1: Click on the discover icon.
(b) Step 2: Click on the first result.
(c) Step 3: Click on the play button.
Figure 7: Multi step grounding case 2: "Open the Mindfulness app, I would like to have a personalized guided meditation to help me be productive throughout the day."